Find symbolic representation for f power -1(x)
f(x)=sq rt x+6

Answer by stanbon(75887)

Find symbolic representation for f power -1(x)
f(x)=sqrt(x+6)
-----------------------
1st: Interchange x and y to get:
x = sqrt(y+6)
--------
2nd: Solve for "y":
Square both sides to get:
y+6 = x^2
----
Then y = x^2 - 6
=====
f^-1(x) = x^2-6
====================
Cheers,
Stan H.
===================